1.The following financial market data is given for an economy (in rupees):

Currency =1000; Reserves =200; Deposits =2000.

a. Calculate size of money multiplier

b. Monetary Base

c. Money Supply


2.State whether the following statements are TRUE or FALSE. Give reason(s) in support of your answer.

(2*1)

i. Higher the marginal propensity to consume, higher is the size of multiplier ii. If investment is very sensitive to interest rate, then we have a flat IS curve.


1
Expert's answer
2021-07-22T13:17:21-0400

 "reserve ratio=reserve\/deposit"

200/2000=0.1

"Money multiplier=1\/Reserve ratio"

=  1/0.1

= 10


b)

"Monetary base=currency + reserves"

= 1000 + 200

= 1200


c)

"money supply=change in reserves *change in money multiplier"

= 10*200

= 2000.


2.


i.) True because the size of the multiplier depends on the household decision to spend, hence the higher the MPC,the higher the size of the multiplier.

ii.)True because Its slope depends on the saving function and investment function.